# Relativistic Mott transition, high-order van Hove singularity, and mean-field phase diagram of twisted double bilayer WSe2

### v1

@@ Abstract
Recent experiments on twisted double bilayer tungsten diselenide have demonstrated that moiré semiconductors can be used to realize a relativistic Mott transition, i.e., a quantum phase transition from a Dirac semimetal to a correlated insulating state, by twist-angle tuning. In addition, signatures of van Hove singularities were observed in the material’s moiré valence bands, suggesting further potential for the emergence of strongly-correlated states in this moiré semiconductor. Based on a Bistritzer-MacDonald-type continuum model, we provide a detailed analysis of the twist-angle dependence of the system’s moiré valence band structure, focusing on both, the evolution of the Dirac excitations and the Fermi-surface structure with its Lifshitz transitions across the van Hove fillings. We exhibit that the twist angle can be used to band engineer a high-order van Hove singularity with power-law exponent 1/4-1/4 in the density of states, which can be accessed by gate tuning of the hole filling. We then study the magnetic phase diagram of an effective Hubbard model for twisted double bilayer tungsten diselenide on the effective moiré honeycomb superlattice with tight-binding parameters fitted to the two topmost bands of the continuum model. To that end, we employ a self-consistent Hartree-Fock mean-field approach in real space. Fixing the angle-dependent Hubbard interaction based on the experimental findings, we explore a broad parameter range of twist angle, filling, and temperature. We find a rich variety of magnetic states that we expect to be accessible in future experiments by twist or gate tuning, including, e.g., a non-coplanar spin-density wave with non-zero spin chirality and a half-metallic uniaxial spin-density wave.
@@ I Introduction
In this work, we provide an extended analysis of the moiré bands and Fermi-surface structure of tdbWSe2, and perform unrestricted Hartree-Fock mean-field calculations to determine the dominant order for the full range of experimentally explored twist angles and fillings. To that end, we start from the continuum-model approach put forward in Ref. [13] and finely resolve the band structure for $\theta\in[2.0^{\circ},4.0^{\circ}]$. One key finding is that each saddle point in the energy dispersion of the second-to-highest moiré band splits into two, when increasing the twist angle beyond $\theta_{c}\sim 3.58^{\circ}$. Right at $\theta_{c}$, the corresponding VHS is then of high order, featuring a power-law divergent DOS instead of the logarithmic one at a conventional saddle point point.
@@ II.1 Continuum model
Due to the translational symmetry, the single-particle Hamiltonian in Eq. (1) can be directly diagonalized by a discrete Fourier transformation. We take into account all Fourier modes within a radius $4G=16\pi^{2}/\sqrt{3}a_{M}$, where $G=|\vec{G}_{i}|$, and we have explicitly verified that the inclusion of more modes leads to negligible quantitative improvement.
@@ III.2 Hubbard interaction
In order to estimate the angle-dependent strength of the on-site Hubbard interaction $U(\theta)$, we employ an approximation scheme put forward in Ref. [18]. To that end, first note that the maxima of the moiré potential associated with the four layers in Eq. (1) coincide to a good approximation (due to the small phase shift of $\phi=-0.17$) and are located on the underlying triangular lattice. Wannier orbitals will then be localized around these maxima, and the on-site interaction can then be estimated by $U\sim e^{2}/4\pi\epsilon a_{W}$}, where $a_{W}$ is the spread of the Wannier function. In the vicinity of the maxima, the potential can be approximated quadratically, yielding $a_{W}\propto\sqrt{a_{M}}$ [18]. Hence, $a_{W}\propto 1/\sqrt{\theta}$, and the on-site Coulomb energy scales as $U\propto\sqrt{\theta}$. The proportionality constant can be fixed by providing the interaction at a reference angle $\theta_{0}$, yielding
@@ V Relativistic Mott transition
In Ref. [12], a transition from a semimetal to an insulator has been observed in tdbWSe2 at $\theta_{c}=2.7^{\circ}$. We use this to fix the interaction strength of our effective Hubbard model, cf. Eq. (8), within the Hartree-Fock approach. The corresponding hopping parameters are already listed in Fig. 1. To proceed, we first determine the value $U_{c}$ as a function of $\theta$, where a transition from a semimetallic to an insulator occurs at Dirac filling. We then set $U(\theta_{0})=U_{c}(\theta_{0})$, where $\theta_{0}=\theta_{c}=2.7^{\circ}$, which fixes the Hubbard interaction for all angles $\theta$, cf. Eq. (9). Compatibility with the experimentally observed transition is then ensured by construction. Our half-filled Hubbard model transitions into a Néel antiferromagnet for $U>U_{c}$.
In order to reliably determine the critical interaction strength $U_{c}$ at which such a phase transition occurs in the thermodynamic limit, we consider the correlation ratio [52]
$R(N,U)=1-\frac{S_{\mathrm{AFM}}(\vec{q}+\delta\vec{q})}{S_{\mathrm{AFM}}(\vec{q})}\,$, (14)
where $\delta\vec{q}$ is a nearest-neighbor reciprocal lattice vector and $S_{\mathrm{AFM}}=S_{\mathrm{AFM}}(\vec{q})$ the static antiferromagnetic structure factor
@@ VI Hartree-Fock phase diagram
We can compare our results to the paradigmatic honeycomb-lattice Hubbard model with nearest-neighbor hopping, only. A detailed Hartree-Fock mean-field study of the latter was recently put forward in Ref. [49] as a benchmark for the investigation of the magnetic states of Bernal bilayer graphene. We find our Hartree-Fock implementation to be consistent with the results presented in Ref. [49]. In particular, for smaller angles $\theta\sim 2^{\circ}$, where the hopping ratios $t_{n>1}/t_{1}$ are systematically smaller than for larger angles and the $U/t_{1}$ is between 3 and 4, we find that the magnetic orders agree very well with the magnetic orders found in Ref. [49]. Minor remaining differences can be explained by the tiny deviations of the free energies of different magnetic states that may induce a preference of one at the expense of another if the hoppings are slightly different.
In summary, our Hartree-Fock mean-field calculations provide robust statements about the presence or absence of order and the accompanied most-dominant momentum transfer. However, the predictions of the precise nature of the order have to be taken with caution. Nevertheless, we interpret the close competition of various different magnetic orders in the phase diagram of the tdbWSe2 model as a characteristic physical feature of the material.
@@
VII Conclusion
@@
In agreement with previous studies, we find that the insulator at half filling is a Néel antiferromagnet, suggesting that the critical point belongs to the Gross-Neveu-Heisenberg universality class [13, 14]. Away from half filling, we then identified various magnetic orders that fall into three main categories: single-mode stripe order, two orthogonal tripes, and three modes with the same amplitude. We find two regions close to the van Hove singularities of the upper and lower bands, where the magnetization is enhanced for twist angles  $\lesssim 2.7^{\circ}$. In this region, the momentum transfer of the dominant spin mode lies at or close to the $m,m^{\prime},m^{\prime\prime}$ points, and the order falls into the category of three modes with the same amplitudes. This allows for the formation of a non-coplanar SDW with non-zero spin chirality [66, 67] or a uniaxial SDW corresponding to a half-metal state [68]. At the high-order van Hove singularity, however, we did not find any enhancement of magnetic order, which is most likely due to an insufficient interaction strength. On the other hand, the high-order van hove singularity may be identified experimentally by its characteristic signal in scanning tunneling spectroscopy [23].
@@
A natural extension of the analysis provided in this work is a fermionic functional renormalization group [69, 70, 71] study of the effective honeycomb-lattice Hubbard model for tdbWSe2. Within the functional renormalization group approach, all electronic many-body instabilities are treated on equal footing, which facilitates a resolution of the competing electronic correlations. Importantly, this approach can reveal unconventional superconductivity in a largely unbiased way. We aim to report on this in a future study. We finally note that the functional renormalization group can also be formulated in a way that explicitly includes order-parameter fluctuations [72]. This can be used to study the quantum critical fan of the Gross-Neveu-Heisenberg universality class in more detail [73]. In conclusion, tdbWSe2 provides an experimentally accessible platform where these theoretical predictions can be tested in a controlled way, opening a pathway towards further aspects of quantum simulation of Dirac fermions [74] and beyond.
@@ Acknowledgements
We thank Aiman Al-Eryani, Jan Biedermann, Laura Classen, Lukas Janssen, Frank Lechermann, Kin Fai Mak, Jonas Profe, and Robin Scholle for insightful discussions and useful correspondence. BH and MMS are supported by the Mercator Research Center Ruhr under Project No. Ko-2022-0012. MMS acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 277146847 (SFB 1238, project C02) and Project No. 452976698 (Heisenberg program).

### v2

# Relativistic Mott transition and high-order van Hove singularity in twisted double bilayer WSe2: mean-field and functional renormalization group study

@@ Abstract
Experiments on twisted double bilayer tungsten diselenide have demonstrated that moir’e semiconductors can realize a relativistic Mott transition, i.e., a quantum phase transition from a Dirac semimetal to a correlated insulating state, by twist-angle tuning. In addition, signatures of van Hove singularities were observed in the material’s moir’e valence bands, suggesting further potential for the emergence of strongly-correlated states. Based on a continuum model, we provide a detailed analysis of the twist-angle dependence of the system’s band structure, focusing on the evolution of the Dirac excitations and the Fermi-surface structure with its Lifshitz transitions across the van Hove fillings. We exhibit that the twist angle can be used to band engineer a high-order van Hove singularity, which can be accessed by gate tuning. We then study the magnetic phase diagram of an effective Hubbard model for twisted double bilayer tungsten diselenide on the effective honeycomb superlattice with tight-binding parameters fitted to the two topmost bands of the continuum model. To that end, we employ a self-consistent Hartree-Fock mean-field approach in real space. Fixing the angle-dependent Hubbard interaction based on the experimental findings, we explore a broad parameter range of twist angle, filling, and temperature. We find a rich variety of magnetic states that we expect to be accessible in future experiments, including, e.g., a non-coplanar spin-density wave with non-zero spin chirality and a half-metallic uniaxial spin-density wave. Finally, we employ a functional renormalization group approach to also study the competition between density-wave and superconducting instabilities. For twist angles of $\theta=2.0^{\circ},2.5^{\circ}$, as well as $\theta\approx 3.5^{\circ}$ – where the high-order van Hove-singularity is found – we find clear indications for unconventional superconductivity.
@@ I Introduction
In this work, we provide an extended analysis of the moiré bands and Fermi-surface structure of tdbWSe2, and perform unrestricted Hartree-Fock mean-field and functional renormalization group calculations to determine the dominant order for the full range of experimentally explored twist angles and fillings. To that end, we start from the continuum-model approach put forward in Ref. [13] and finely resolve the band structure for $\theta\in[2.0^{\circ},4.0^{\circ}]$. One key finding is that each saddle point in the energy dispersion of the second-to-highest moiré band splits into two, when increasing the twist angle beyond $\theta_{c}\sim 3.58^{\circ}$. Right at $\theta_{c}$, the corresponding VHS is then of high order, featuring a power-law divergent DOS instead of the logarithmic one at a conventional saddle point point.
@@
Finally, we employ a functional renormalization group (fRG) approach to explore the competition of density-wave instabilities, including the tendencies towards magnetic order, and superconducting instabilities.
@@ II.1 Continuum model
Due to the translational symmetry, the single-particle Hamiltonian in Eq. (1) can be directly diagonalized by a discrete Fourier transformation. We take into account all Fourier modes within a radius $4G=16\pi/\sqrt{3}a_{M}$, where $G=|\vec{G}_{i}|$, and we have explicitly verified that the inclusion of more modes leads to negligible quantitative improvement.
@@ III.2 Hubbard interaction
In order to estimate the angle-dependent strength of the on-site Hubbard interaction $U(\theta)$, we employ an approximation scheme put forward in Ref. [18]. To that end, first note that the maxima of the moiré potential associated with the four layers in Eq. (1) coincide to a good approximation (due to the small phase shift of $\phi=-0.17$) and are located on the underlying triangular lattice. Wannier orbitals will then be localized around these maxima, and the on-site interaction can then be estimated by $U\sim e^{2}/4\pi\epsilon a_{W}$}, where $a_{W}$ is the spread of the Wannier function. In the vicinity of the maxima, the potential can be approximated quadratically, yielding $a_{W}\propto\sqrt{a_{M}}$ [18]. Hence, $a_{W}\propto 1/\sqrt{\theta}$, and the on-site Coulomb energy scales as $U\propto\sqrt{\theta}$.
The proportionality constant can be fixed by providing the interaction at a reference angle $\theta_{0}$, yielding
@@ V Relativistic Mott transition
In Ref. [12], a transition from a semimetal to an insulator has been observed in tdbWSe2 at $\theta_{c}=2.7^{\circ}$. We use this to fix the interaction strength of our effective Hubbard model, cf. Eq. (8), within the Hartree-Fock approach. The corresponding hopping parameters are already listed in Fig. 1.
To proceed, we first determine the value $U_{c}$ as a function of $\theta$, where a transition from a semimetallic to an insulator occurs at Dirac filling. We then set $U(\theta_{0})=U_{c}(\theta_{0})$, where $\theta_{0}=\theta_{c}=2.7^{\circ}$, which fixes the Hubbard interaction for all angles $\theta$, cf. Eq. (9). Compatibility with the experimentally observed transition is then ensured by construction. Our half-filled Hubbard model transitions into a Néel antiferromagnet for $U>U_{c}$.
In order to reliably determine the critical interaction strength $U_{c}$ at which such a phase transition occurs in the thermodynamic limit, we consider the correlation ratio [52, 53, 54]
$R(N,U)=1-\frac{S_{\mathrm{AFM}}(\vec{q}+\delta\vec{q})}{S_{\mathrm{AFM}}(\vec{q})}\,.$ (14)
Here, the vector $\delta\vec{q}$ is a nearest-neighbor reciprocal lattice vector and $S_{\mathrm{AFM}}=S_{\mathrm{AFM}}(\vec{q})$ is the static antiferromagnetic structure factor, reading
@@ VI Hartree-Fock phase diagram
We can compare our results to the paradigmatic honeycomb-lattice Hubbard model with nearest-neighbor hopping, only. A detailed Hartree-Fock mean-field study was recently put forward in Ref. [49] as a benchmark for the investigation of Bernal bilayer graphene. We find our implementation to be consistent with the results presented there. In particular, for smaller angles $\theta\sim 2^{\circ}$, where the hopping ratios $t_{n>1}/t_{1}$ are systematically smaller than for larger angles and the $U/t_{1}$ is between 3 and 4, we find that the magnetic orders agree very well with the ones in Ref. [49]. Minor differences can be explained by the tiny deviations of the free energies of different magnetic states that may induce a preference of one at the expense of another if the hoppings are slightly different.
In summary, our Hartree-Fock calculations provide robust statements about the presence or absence of order and the accompanied most-dominant momentum transfer. However, the predictions of the precise nature of the order have to be taken with caution. We interpret the close competition of various different magnetic orders in the phase diagram of the tdbWSe2 model as a characteristic feature of the material.
@@
VII Functional renormalization group calculations
VII.1 Functional renormalization group method
The Hartree-Fock mean-field study from the previous chapter exhibited the system’s strong tendency towards the formation of magnetic order for smaller twist angles. Notably, spin fluctuations may also provide the pairing glue for the formation of (unconventional) superconductivity, which is, however, not included in the Hartree-Fock approach. To study competing correlated phases, including superconductivity, in the effective honeycomb-lattice Hubbard model for tdbWSe2, we therefore employ a fermionic fRG approach, which identifies the leading many-body instabilities including various types of density-wave and superconducting instabilities on equal footing [70, 71, 72, 73].
In our approximation, we exclusively focus on the renormalization group behavior of the normal two-particle interaction vertex $\Gamma^{(4)}$, which is spin-rotation invariant in our model and depends on three independent wavevectors. Schematically, the fRG introduces a scale parameter $\Lambda$ to interpolate smoothly between the bare interactions at the ultraviolet (UV) scale $\Lambda=\Lambda_{\mathrm{UV}}$ and the fully dressed ones at $\Lambda=0$. Our regularization scheme is based on a sharp cutoff in frequencies [74, 71, 75] so that the vertex at scale $\Lambda$ dresses the bare interaction with all fluctuations of energy above that scale. In our approximation, we neglect self-energy corrections, frequency dependencies, and higher-order vertices with more than four fermionic fields. This truncation scheme has been frequently employed in the context of strongly-correlated electron systems to determine Fermi liquid instabilities in an unbiased way [70, 71, 76, 72, 73, 77].
A flow to strong coupling of $\Gamma^{(4)}$ at finite $\Lambda=\Lambda_{c}$ signals an instability of the normal state towards a symmetry broken phase. Using the effective vertex at the critical scale $\Lambda_{c}$, we classify the instability either as a transition towards a spin/charge-density-wave state (DW) or a superconducting state (SC). For the present system, we have employed a recent implementation of the truncated-unity FRG [78, 79, 80, 81, 82], which is based on a decomposition of the interaction vertex in terms of three scattering channels, i.e, $P$, $D$ and $C$, each of which depends most strongly on the Mandelstam variables for pairing $q_{P}$, direct $q_{D}$ and crossed $q_{C}$ channels, respectively [74]. The remaining momentum dependencies are weak and may be expanded into a truncated set of form factors that respect the symmetry of the problem at a minimal loss of accuracy [74, 78, 79].
For our practical fRG calculations, we made use of the divERGe library [81, 82], employing the TUFRG backend with a wavevector resolution of the Mandelstam variables with $N_{q}=50^{2}$ points in the Brillouin zone. The wave-vector resolution in the loop integrations requires higher accuracy and we therefore choose a resolution of $N_{l}=2500^{2}$. The truncated form-factor expansion includes all form-factor bonds connecting sites at distance equal to or smaller than $1a_{M}$ from either of the two sites of a reference unit cell, making up a set of eight bond form factors.
Figure 6: fRG results for competing correlated phases. We show results of our fRG calculations for the critical scale $\Lambda_{c}$ as a function of the filling $\nu$ and for twist angles $\theta=2.0^{\circ},2.5^{\circ},3.51^{\circ}$. The latter angle includes the HOVHS near filling $\nu\approx 2.2$. Closed circles indicate spin DW instabilities, whereas triangles represent SC instabilities. The color of the symbols indicates the ordering vector associated to the instability. All superconducting instabilities are found to be in the $E_{2}$ irreducible representation of the point group of the honeycomb lattice.
VII.2 Competing correlated phases from fRG
To set up the numerical evaluation of the fRG for tdbWSe2, we start by gauging the interaction values $U(\theta)$ with the experimental observation of the relativistic Mott transition occurring at the angle $\theta_{0}=2.7^{\circ}$ by using Eq. (9) to set $U(\theta_{0})$. This is necessary because the fRG includes additional fluctuations beyond mean-field, which also tentatively mitigate the overestimation of ordering within mean-field approaches. Indeed, we find the instability signaling this transition occurring at $U^{\mathrm{fRG}}(\theta_{0})\approx 3.4t_{1}(\theta_{0})$, which is closer to the numerical results from QMC [41, 43, 45] as compared to the mean-field value.
Having fixed the Hubbard interaction in this way, we now choose three representative twist angles: (i) $\theta=2.0^{\circ}$ features the approximately-nested Fermi surface at the hole-doped van Hove filling and the Hubbard interaction is already relatively strong, i.e., $U(\theta=2.0^{\circ})\approx 5.5t_{1}$, (ii) $\theta=3.51^{\circ}$ exhibits a HOVHS at hole-doping albeit at smaller Hubbard interaction, i.e., $U(\theta=3.51^{\circ})\approx 2.2t_{1}$, and (iii) $\theta=2.5^{\circ}$ represents an intermediate scenario where $U(\theta=2.5^{\circ})\approx 3.9t_{1}$. Adjusting the chemical potential, we again performed parameter scans in the filling between the VHS in the two honeycomb bands and our results are shown in Fig. 6.
For $\theta=2.0^{\circ}$ and $\theta=2.5^{\circ}$ and at Dirac filling, we find a divergence of the particle-hole exchange vertex $C$ at ordering vector $\gamma$. Diagonalizing the vertex $C(q=\gamma)$ in form-factor space, we find the leading eigenvector to have opposite spin projection on the two sublattices, indicating an insulating Néel ground state, in agreement with our Hartree-Fock study in Sec. VI. As we dope away slightly, we find an instability towards incommensurate spin DWs with momentum transfers that are also in reasonable agreement with the ones in the magnetic orders described in the mean-field analysis, cf. the middle panels of Fig. 4. The fRG also independently provides further evidence for the tendencey towards magnetic order with momentum transfer $m$ near the van Hove fillings, wherever this is the leading instability.
For fillings between the van Hove and the Dirac fillings we had already found in the Hartree-Fock study that the tendency towards magnetic order is somewhat weaker. This can be seen, e.g., in the magnetization in Fig. 4, which becomes smaller or even vanishes in this region for $\theta=2.5^{\circ}$. In the fRG analysis, we find that for these fillings superconducting instabilities develop, instead, see Fig. 6. Diagonalization of the pairing vertex $P(q)$ in form-factor space at $q=\gamma$, identifies two degenerate leading eigenvectors that span the $E_{2}$ irreducible representation of the point group $C_{6v}$. This suggests that the superconducting state is composed of a superposition of the $d_{x^{2}-y^{2}}$- and $d_{xy}$-wave form factors. It can be further expected that a superposition of the type $d_{x^{2}-y^{2}}\pm id_{xy}$, i.e., a chiral $d$-wave superconducting state, which opens a full gap, is energetically favored [83].
Remarkably, and in contrast to our observations in the Hartree-Fock analysis, we also find many-body instabilities near the HOVHS for $\theta=3.51^{\circ}$. This may be partly attributed to the slightly larger Hubbard interactions that are required to fit the experimental findings on the relativistic Mott transition. More specifically, right at the high-order van Hove filling, we find an instability in the $C$ channel at $q=\gamma$. Further analysis of $C(q=\gamma)$ reveals the leading eigenvector to have equal spin projection on the two sublattices, indicating a ferromagnetic ground state in agreement with Ref. [28]. Moving the filling slightly away from the HOVHS, we find again $E_{2}$ superconductivity, however, with steeply decaying $\Lambda_{c}$. In the remaining region of fillings no further instabilities are found for $\theta=3.51$.
VIII Conclusion
@@
In agreement with previous studies, we find that the insulator at half filling is a Néel antiferromagnet, suggesting that the critical point belongs to the Gross-Neveu-Heisenberg universality class [13, 14]. Away from half filling, we then identified various magnetic orders that fall into three main categories: single-mode stripe order, two orthogonal tripes, and three modes with the same amplitude. We find two regions close to the van Hove singularities of the upper and lower bands, where the magnetization is enhanced for twist angles  $\lesssim 2.7^{\circ}$. In this region, the momentum transfer of the dominant spin mode lies at or close to the $m,m^{\prime},m^{\prime\prime}$ points, and the order falls into the category of three modes with the same amplitudes. This allows for the formation of a non-coplanar SDW with non-zero spin chirality [67, 68] or a uniaxial SDW corresponding to a half-metal state [69]. At the high-order van Hove singularity, however, we did not find any enhancement of magnetic order within our Hartree-Fock mean-field approach, which is most likely due to an insufficient interaction strength. On the other hand, the high-order van Hove singularity may be identified experimentally by its characteristic signal in scanning tunneling spectroscopy [23].
@@
As a natural extension of the Hartree-Fock analysis we carried out a fermionic functional renormalization group study of the effective honeycomb-lattice Hubbard model for tdbWSe2. Within the functional renormalization group approach, all electronic many-body instabilities are treated on equal footing, which facilitates a resolution of the competing electronic correlations. Importantly, this approach revealed unconventional superconductivity with a pairing gap in the $E_{2}$ irrep of the point group, suggesting chiral $d$-wave superconductivity. We furthermore find an enhancement of many-body instabilities close to the HOVHS at $\theta\approx 3.51^{\circ}$, which is in contrast to our findings from Hartree-Fock mean-field theory. In particular, we identify a ferromagnetic ground state right at the high-order van Hove filling, and $E_{2}$ superconductivity slightly away from it.
We finally note that the functional renormalization group can also be formulated in a way that explicitly includes order-parameter fluctuations [73]. This can be used to study the quantum critical fan of the Gross-Neveu-Heisenberg universality class in more detail [84]. In conclusion, tdbWSe2 provides an experimentally accessible platform where these theoretical predictions can be tested in a controlled way, opening a pathway towards further aspects of quantum simulation of Dirac fermions [85] and beyond.
@@ Acknowledgements
We thank Jan Biedermann, Laura Classen, Lukas Janssen, Frank Lechermann, Kin Fai Mak, and Robin Scholle for insightful discussions and useful correspondence. We especially thank Jonas Profe for helpful discussions and practical advice concerning the divERGe library. BH and MMS are supported by the Mercator Research Center Ruhr under Project No. Ko-2022-0012. MMS acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 277146847 (SFB 1238, project C02) and Project No. 452976698 (Heisenberg program).